# Chatbot Arena Meets Nuggets: Towards Explanations and Diagnostics in the Evaluation of LLM Responses

### v1

@@ Abstract
Battles, or side-by-side comparisons in so-called arenas that elicit human preferences, have emerged as a popular approach to assessing the output quality of LLMs. Recently, this idea has been extended to retrieval-augmented generation (RAG) systems. While undoubtedly representing an advance in evaluation, battles have at least two drawbacks, particularly in the context of complex information-seeking queries: they are neither explanatory nor diagnostic. Recently, the nugget evaluation methodology has emerged as a promising approach to evaluate the quality of RAG answers. Nuggets decompose long-form LLM-generated answers into atomic facts, highlighting important pieces of information necessary in a “good” response. In this work, we apply our AutoNuggetizer framework to analyze data from roughly 7K Search Arena battles provided by LMArena in a fully automatic manner. Our results show a significant correlation between nugget scores and human preferences, showcasing promise in our approach to explainable and diagnostic system evaluations.
Chatbot Arena Meets Nuggets: Towards Explanations and Diagnostics
in the Evaluation of LLM Responses
Sahel Sharifymoghaddamthanks: Equal Contribution, Shivani Upadhyay11footnotemark: 1, Nandan Thakur11footnotemark: 1, Ronak Pradeep, Jimmy Lin David R. Cheriton School of Computer Science, University of Waterloo, Canada {sahel.sharifymoghaddam, sjupadhyay, nandan.thakur rpradeep, jimmylin@}uwaterloo.ca
@@ 1 Introduction
The notion of “battles”, or side-by-side comparisons of responses from large language models (LLMs), has become a popular method for evaluating the quality of LLMs Zheng et al. (2023); Chiang et al. (2024). In the “arena” setup, users are shown two LLM outputs and asked to indicate which one they prefer. This approach was popularized by LMSYS through MT-Bench Zheng et al. (2023) and later expanded into the Chatbot Arena Chiang et al. (2024). The popularity of these arenas has made them a key marketing tool when launching new LLMs from companies such as Google, OpenAI, and Meta, who regularly tout leaderboard rankings on Chatbot Arena in model releases. Inspired by this increased popularity, arena-based evaluations have been extended to a variety of domains, including AI agents Yekollu et al. (2024), vision and image generation Lu et al. (2024); Jiang et al. (2024), and multilingual generation Thakur et al. (2025a).
@@
Most recently, the idea of battles was extended to search-augmented LLMs in the Search Arena Miroyan et al. (2025). Unlike the original setup, which focused on “closed-book” generation of responses by LLMs directly, Search Arena evaluates systems that apply retrieval-augmented generation (RAG) to retrieve relevant source documents, which are then used by LLMs to generate long-form answers with citations Pradeep et al. (2025a); Han et al. (2024). While such side-by-side comparisons enable the evaluation of search-augmented LLM-based systems at scale, we see them having at least two drawbacks: they are neither explanatory nor diagnostic, especially in scenarios where determining the better answer is not straightforward. It would be desirable for an evaluation to (at least attempt to) explain why a user might have preferred one response over another. Furthermore, we argue that evaluations should be diagnostic in providing actionable guidance on how to improve systems.
@@
We hypothesize that the recently introduced nugget evaluation methodology Pradeep et al. (2024, 2025b) can be adapted to potentially address these two drawbacks for complex information-seeking queries. The basic idea is to assess answer quality in terms of the recall of information nuggets, or atomic facts, that should be present in good responses. In our previous work, we have shown that this can be accomplished in a fully automatic manner using LLMs.
@@
Figure 1: An end-to-end example from Search Arena illustrating both nugget generation and assignment. First, GPT4.1 generates nuggets based on the query and the responses from both models. Each nugget is then labeled with an importance level—either “vital” or “okay”. Next, GPT4.1 evaluates whether each model supports each nugget, assigning one of three labels: “support”, “partial support”, or “no support”. Finally, these support judgments are scored and aggregated to determine the overall outcome (the model with the higher score is preferred).
@@
In this paper, we adapt the AutoNuggetizer Pradeep et al. (2024) implementation of the nugget evaluation methodology to analyze recently released public data from Search Arena in a fully automatic manner (see Figure 1). We find that human preferences correlate well with the distribution of nugget scores, conditioned on these preferences, which can be plotted as density functions. Further analyses reveal that these distributions differ significantly from one another, providing strong evidence for the explanatory power of nugget scores in capturing human preferences. From here, nugget score differences provide actionable guidance to improve RAG systems.
@@ 2 Related Work
Originally introduced in the TREC QA Track in 2003 (Voorhees, 2003a, b), the nugget-based evaluation methodology focuses on identifying essential atomic facts—called nuggets—that are relevant to a given question. This methodology was later extended to tasks like summarization and broader conceptions of question answering (Nenkova and Passonneau, 2004; Lin and Demner-Fushman, 2006b; Dang and Lin, 2007; Lin and Zhang, 2007), and researchers have explored automation to improve its scalability (Lin and Demner-Fushman, 2005, 2006a; Pavlu et al., 2012).
The recent emergence of large language models (LLMs) has enabled automated, reliable nugget-based evaluation (Pradeep et al., 2024; Alaofi et al., 2024; Pradeep et al., 2025b; Thakur et al., 2025b; Abbasiantaeb et al., 2025). Several RAG evaluation frameworks—such as FactScore (Min et al., 2023), RUBRIC (Farzi and Dietz, 2024), and others (Arabzadeh and Clarke, 2024; Mayfield et al., 2024)—incorporate the nugget concept, although most of these proposed approaches are either not validated or primarily validated on traditional ad hoc retrieval, and hence their applicability to long-form answers is unclear. We refer readers to Pradeep et al. (2025b) for a more detailed discussion of related work. Here, we use the AutoNuggetizer framework from Pradeep et al. (2024), applying it to side-by-side comparisons of LLM responses within the Search Arena.
@@
Search Arena, introduced by LMArena (Miroyan et al., 2025), is a popular benchmark focused on evaluation of LLMs with access to a search tool. Other notable efforts include the MTEB Arena Hugging Face (2023), which extends the Massive Text Embedding Benchmark (MTEB) framework Muennighoff et al. (2023) to head-to-head evaluation across embedding models, and Ragnarök (Pradeep et al., 2025a), which released a new variant of the MS MARCO collection (V2.1) and offers a framework for evaluation in the TREC 2024 RAG Track.
@@ 3 Experimental Design
The Search Arena, introduced by LMArena in Miroyan et al. (2025), is a crowd-sourced evaluation platform for search-augmented LLM systems, evaluated in terms of side-by-side comparisons that solicit human preferences Chiang et al. (2024). The V1 version of the publicly released dataset111huggingface.co/datasets/lmarena-ai/search-arena-v1-7k contains in total 7K samples for which two RAG focused systems (e.g., Gemini-2.5-Pro-Grounding vs. Perplexity-Sonar-Reasoning-Pro) battle each other. For each battle, a human assessor judges with one of four responses, whether modelA (or modelB) is the winner or a tie where both are responses are good (or bad).
The Search Arena dataset contains both single-turn and multi-turn battles. In this work, we focus exclusively on single-turn battles, evaluating 5,103 instances. We excluded multi-turn battles from our experiments because human votes at the overall battle level do not reliably reflect turn-level (per-query) preferences, which is what AutoNuggetizer evaluates. Search Arena also contains battles for several non-English languages, e.g., Chinese or Russian. Non-English languages collectively account for less than 40% of the dataset, with English comprising the remaining majority. Detailed statistics for single-turn battles used in this work are presented in Section A.1. Queries in Search Arena are diverse; they can be very long (a code snippet), reasoning-intensive, ambiguous, and occasionally vague. We show some examples of queries from the dataset in Section A.2.
Nugget evaluations.
Nugget generation creates atomic facts that highlight the essential information required in a RAG answer. Following Pradeep et al. (2024), we use the AutoNuggetizer tool in the nuggetizer code repository [^2] https://github.com/castorini/nuggetizer to generate and assign information nuggets to model responses. As shown in Figure 1, there are two steps in nugget generation and assignment:
@@
    Nugget generation: For each prompt extracted from the dataset, we construct a request to AutoNuggetizer consisting of a query (the prompt itself) and two documents, which are responses from each model, randomly ordered to avoid positional bias. The tool identifies nuggets that are relevant to the query from the provided LLM responses. Furthermore, each nugget is assigned an importance label: “vital” or “okay”, reflecting its relevance or significance to the query. Following previous guidelines, “vital” nuggets are those that must be present in a “good” answer, while “okay” nuggets are nice to have, but are not absolutely necessary. Nugget importance labeling is run in a separate pass, independent of the actual responses.
@@
    Nugget assignment: Once nuggets and their importance labels are generated (from the previous step), we then use AutoNuggetizer to assign them to model responses, determining whether the nugget is found in the answer. This step categorizes each nugget into “supported”, “partially supported”, or “not supported”. Among the four combinations of evaluation methods—“vital” vs. “all” (vital + okay) and “strict” vs. “non-strict” (full + partial support)—we adopt the “All Score” metric, which achieves the highest recall by counting nuggets of all importance and support levels. We find that the “Strict Vital” metric, which is the primary metric used in the TREC 2024 RAG Track Pradeep et al. (2025b), is too strict for our use case, particularly when only a small number of nuggets are available.
We emphasize that while the AutoNuggetizer framework supports different degrees of manual intervention, here we are running the entire evaluation pipeline end-to-end in an automatic manner.
@@ 4 Experimental Results
All experiments in this paper are conducted using GPT4.1, the latest language model from OpenAI, as the underlying model used by the AutoNuggetizer via Microsoft Azure. Out of the 5,103 single-turn battles in the Search Arena dataset, 51 were excluded from our analysis due to issues such as Azure content filtering, invalid output formats, or other nugget generation failures.
Figure 2: Empirical probability density function (PDF) of nugget score differences ($\textrm{score}_{B}$- $\textrm{score}_{A}$) grouped by human vote category: modelA wins, tie, or modelB wins. A Kernel Density Estimation (KDE) with a bandwidth of 0.5 is fitted separately for each group.
@@
These results appear to support our hypothesis that nugget scores correlate with human preferences. In the case where modelA wins (top row), the distribution skews to the left (negative values), indicating that modelA typically gets higher nugget scores than modelB. Conversely, when modelB wins (bottom row), the distribution skews to the right (positive values), suggesting that modelB generally obtains higher nugget score. When the human indicates a tie (middle row), the distribution peaks around zero, as expected, indicating similar nugget scores between models.
To analyze the statistical differences among these three conditional distributions, we performed pairwise Kolmogorov-Smirnov (K-S) tests. As shown in Figure 3, the K-S statistic values range from 0.211 to 0.324, with $p$-values of $1.3\mathrm{e}^{-25}$ or lower, indicating that all three distributions differ significantly from one another (i.e., we have high confidence that these samples were drawn from different distributions). These findings validate our hypothesis that nugget score differences align with human preferences, reinforcing the potential of nugget-based metrics as reliable evaluators of model quality in head-to-head evaluations.
Figure 4: Confusion matrix comparing human and nugget preferences. Threshold of 0.1 is used to treat the nugget preference as a tie.
Figure 4 presents a confusion matrix that visually compares the distribution of human preferences (rows) in Search Arena against “nugget preferences” (columns). For “nugget preference”, we use a threshold of 0.1, meaning that when the nugget score difference for the two model outputs falls within ±$\pm 0.1$, the comparison is considered a tie. The diagonal cells in the confusion matrix reveal the instances where nugget preferences align with human preferences. Conversely, off-diagonal cells illustrate the types and frequencies of disagreements between the human and nugget scores.
In particular, the nugget-based evaluation prefers modelA in 933 out of 1704 (54.8%) of the battles where modelA wins the battle (first row in Figure 4). Similarly, modelB is preferred in 911 out of 1740 (52.4%) battles where it wins the battle (third row in Figure 4). Lastly, when a tie occurs, roughly similar preferences are assigned to all three choices (second row in Figure 4). We investigate the anti-diagonal corners where nugget and human preferences disagree in the studies below.
(a) Multi-faceted
(b) Ambiguous
(c) Assumptive

### v2

@@ Abstract
Battles, or side-by-side comparisons in so-called arenas that elicit human preferences, have emerged as a popular approach for assessing the output quality of LLMs. Recently, this idea has been extended to retrieval-augmented generation (RAG) systems. While undoubtedly representing an advance in evaluation, battles have at least two drawbacks, particularly in the context of complex information-seeking queries: they are neither explanatory nor diagnostic. Recently, the nugget evaluation methodology has emerged as a promising approach to evaluate the quality of RAG answers. Nuggets decompose long-form LLM-generated answers into atomic facts, highlighting important pieces of information necessary in a “good” response. In this work, we apply our AutoNuggetizer framework to analyze data from roughly 7K Search Arena battles provided by LMArena in a fully automatic manner. Our results show a significant correlation between nugget scores and human preferences, showcasing promise in our approach to explainable and diagnostic system evaluations. All the code necessary to reproduce results in our work is available in https://github.com/castorini/lmsys_nuggetize.
@@ 1 Introduction
The notion of “battles”, or side-by-side comparisons of responses from large language models (LLMs), has become a popular method for evaluating their quality [32, 4]. In the “arena” setup, users are shown two LLM outputs and asked to indicate which one they prefer. This approach was popularized by LMSYS through MT-Bench [32] and later expanded into the Chatbot Arena [4]. The popularity of these arenas has made them a key marketing tool when launching new LLMs from companies such as Google, OpenAI, and Meta, who regularly tout leaderboard rankings on Chatbot Arena in model releases. Inspired by this increased popularity, arena-based evaluations have been extended to a variety of domains, including AI agents [31], vision and image generation [14, 8], and multilingual generation [27].
@@
Most recently, battles were extended to search-augmented LLMs in the Search Arena [18]. Unlike the original setup, which focused on “closed-book” generation of responses by LLMs directly, Search Arena evaluates systems that apply retrieval-augmented generation (RAG) to retrieve relevant source documents, which are then used by LLMs to generate long-form answers with citations [22, 7]. While such side-by-side comparisons enable the evaluation of search-augmented LLM-based systems at scale, we see them having at least two drawbacks: they are neither explanatory nor diagnostic, especially in scenarios where determining the better answer is not straightforward. It would be desirable for an evaluation to (at least attempt to) explain why a user might have preferred one response over another. Furthermore, we argue that evaluations should be diagnostic in providing actionable guidance on how to improve systems.
@@
We hypothesize that the recently introduced nugget evaluation methodology [23, 24] can be adapted to potentially address these two drawbacks for complex information-seeking queries. The basic idea is to assess answer quality in terms of the recall of information nuggets, or atomic facts, that should be present in good responses. In our previous work, we have shown that this can be accomplished in a fully automatic manner using LLMs.
@@
Figure 1: An end-to-end example from Search Arena illustrating both nugget generation and assignment. First, GPT4.1 generates nuggets based on the query, retrieved chunks from URL contents, and the responses from both models. Each nugget is then labeled with an importance level—either “vital” or “okay”. Next, GPT4.1 evaluates whether each model supports each nugget, assigning one of three labels: “support”, “partial support”, or “no support”. Finally, these support judgments are scored and aggregated to determine the overall outcome (the model with the higher score is preferred).
@@
In this paper, we adapt the AutoNuggetizer [23] implementation of the nugget evaluation methodology to analyze recently released public data from Search Arena in a fully automatic manner (see Figure 1). We find that human preferences correlate well with the distribution of nugget scores, conditioned on these preferences, which can be plotted as density functions. Further analyses reveal that these distributions differ significantly from one another, providing strong evidence for the explanatory power of nugget scores in capturing human preferences. From here, nugget score differences provide actionable guidance to improve RAG systems.
@@ 2 Related Work
Originally introduced in the TREC QA Track in 2003 [29, 30], the nugget-based evaluation methodology focuses on identifying essential atomic facts—called nuggets—that are relevant to a given question. This methodology was later extended to tasks like summarization and broader conceptions of question answering [20, 11, 5, 13], and researchers have explored automation to improve its scalability [9, 10, 21].
The recent emergence of large language models (LLMs) has enabled automated, reliable nugget-based evaluation [23, 2, 24, 28, 1]. Several RAG evaluation frameworks—such as FactScore [17], RUBRIC [6], and others [3, 15]—incorporate the nugget concept, although most of these proposed approaches are either not validated or primarily validated on traditional ad hoc retrieval, and hence their applicability to long-form answers is unclear. We refer readers to Pradeep et al. [24] for a more detailed discussion of related work. Here, we use the AutoNuggetizer framework from Pradeep et al. [23], and apply it to the side-by-side comparisons of LLM responses within the Search Arena.
@@
The Search Arena, introduced by LMArena [18], is a recently introduced (April 14, 2025) and popular benchmark evaluating LLMs with access to a search tool. Other notable efforts include the MTEB Arena [16], which extends the Massive Text Embedding Benchmark (MTEB) framework [19] to head-to-head evaluation across embedding models, and Ragnarök [22], which released a new variant of the MS MARCO collection (V2.1) and offers a framework for evaluation in the TREC 2024 RAG Track.
@@ 3 Experimental Design
The Search Arena, introduced by LMArena in Miroyan et al. [18], is a crowd-sourced evaluation platform for search-augmented LLM systems, evaluated in terms of side-by-side comparisons that solicit human preferences [4]. The V1 version of the publicly released dataset [^1] https://huggingface.co/datasets/lmarena-ai/search-arena-v1-7k contains in total 7K samples for which two RAG-focused systems denoted as modelA and modelB (e.g., Gemini-2.5-Pro-Grounding vs. Perplexity-Sonar-Reasoning-Pro) battle each other. For each battle, a human assessor judges with one of four responses, whether modelA (or modelB) is the winner or a tie where both are responses are good (or bad). Each battle also includes search result URLs used by the search-augmented LLM systems for response generation when available—approximately 6.7K out of 7K battles. In total, the dataset contains around 80K unique URLs.
The Search Arena dataset comprises both single-turn and multi-turn battles. In this work, we focus exclusively on single-turn battles, where the system provides a single response back to the user, evaluating a total of 5,103 instances. We exclude multi-turn battles from our experiments because human votes at the overall battle level do not reliably reflect turn-level (per-query) preferences, which is what AutoNuggetizer evaluates.
Search Arena also contains battles for several non-English languages, e.g., Chinese or Russian. Non-English languages collectively account for less than 40% of the dataset, with English comprising the remaining majority. Detailed statistics for single-turn battles used in this work are presented in Section A.1. Queries in Search Arena vary widely, ranging from long code snippets to prompts that demand complex reasoning or exhibit ambiguity and vagueness. We show a few examples of queries from the Search Arena dataset in Section A.2.
Corpus generation.
To evaluate LLM responses in the absence of ground-truth answers, we use the URLs provided in the dataset, collected from each system response as relevant sources of information. We begin by constructing a corpus from the 47K unique URLs associated with single-turn battles. This process involves downloading the contents of each URL, extracting the main textual content, and segmenting the text into chunks of ten sentences with an overlap of two sentences, using spaCy’s  [^2] https://spacy.io/ xx_sent_ud_sm model.
Once the corpus is prepared, we encode the chunks and the query prompts utilizing the BAAI/bge-m3 model. We then perform dense retrieval by computing cosine similarity between the chunk and query embeddings, implemented via Pyserini’s FAISS indexing and search [12]. With this setting, we retrieve the top 50 most relevant chunks for each query. Notably, both the chunking and encoding models support multilingual corpora, ensuring robust language coverage across the dataset.
Figure 2: Empirical probability density function (PDF) of nugget score differences ($\textrm{score}_{B}$- $\textrm{score}_{A}$) grouped by human preference category: modelA wins, tie, or modelB wins. A Kernel Density Estimation (KDE) with a bandwidth of 0.5 is fitted separately for each group.
Nugget evaluation.
Nugget generation creates atomic facts that highlight the essential information required in a RAG answer, and assignment categorizes their support level for the model response. Following Pradeep et al. [23], we use the AutoNuggetizer tool in the nuggetizer code repository [^3] https://github.com/castorini/nuggetizer to generate and assign information nuggets to model responses. As shown in Figure 1, there are two steps in nugget generation and assignment:
@@
    Nugget generation: For each prompt extracted from the dataset, we construct a request to AutoNuggetizer that includes the query (i.e., the prompt itself), along with relevant chunks retrieved from our created corpus, ordered by relevance. In addition to these chunks, we include the responses from each model, inserted in random order to mitigate the positional bias. We include model responses for two key reasons. First, approximately 5% of the battles do not contain any URLs. Second, even when URLs are present, about 16% of them yield 100 bytes or less of content after scraping–typically due to issues such as cookie or JavaScript requirements, invalid or expired links, geo-blocking, and similar obstacles. These cases make the LLM responses a valuable fallback source of information for nugget generation.
    The AutoNuggetizer tool then processes the request and identifies nuggets that are relevant to the query from the retrieved chunks and the provided LLM responses. Furthermore, each nugget is assigned an importance label: “vital” or “okay”, reflecting its relevance or significance to the input query. Following previous guidelines, “vital” nuggets must be present in a “good” answer, while “okay” nuggets are good to have, but are not necessary. Nugget importance labeling is run in a separate pass, independent of the actual responses.
@@
    Nugget assignment: Once nuggets and their importance labels are generated (from the previous step), we use AutoNuggetizer to assign them to model responses, determining whether each nugget is supported in the answer. This step categorizes each nugget into “supported”, “partially supported”, or “not supported”. Among the four combinations of evaluation methods—“vital” vs. “all” (vital + okay) and “strict” vs. “non-strict” (full + partial support)—we adopt the “All Score” metric, which achieves the highest recall by counting nuggets of all importance and support levels. We find that the “Strict Vital” metric, which is the primary metric used in the TREC 2024 RAG Track [24], is too strict for our use case, particularly when only a small number of nuggets are available.
We emphasize that while the AutoNuggetizer framework supports different degrees of manual intervention, here we are running the entire evaluation pipeline end-to-end automatically.
@@ 4 Experimental Results
All experiments in this paper are conducted using GPT4.1, the latest language model from OpenAI with a knowledge cutoff of June 2024, as the underlying model used by the AutoNuggetizer via Microsoft Azure. Out of the 5,103 single-turn battles in the Search Arena dataset, 5 were excluded from our analysis due to issues such as Azure content filtering, invalid output formats, or other nugget generation failures. On average, each single-turn battle full evaluation (comprising both nuggetization and assignment) requires approximately 2–3 seconds when executed using the Azure OpenAI API. The mean number of nuggets generated per battle is $\sim$12.5.
These results appear to support our hypothesis that nugget scores correlate with human preferences. In the case where modelA wins (top row), the distribution skews to the left (negative values), indicating that modelA typically gets higher nugget scores than modelB. Conversely, when modelB wins (bottom row), the distribution skews to the right (positive values), suggesting that modelB generally obtains a higher nugget score. When the human indicates a tie (middle row), the distribution peaks around zero, as expected, indicating similar nugget scores between models.
To analyze the statistical differences among these three conditional distributions, we performed pairwise Kolmogorov-Smirnov (K-S) tests. As shown in Figure 3, the K-S statistic values range from 0.205 to 0.313, with $p$-values of $1.2\mathrm{e}^{-24}$ or lower, indicating that all three distributions differ significantly from one another (i.e., we have high confidence that these samples were drawn from different distributions). These findings validate our hypothesis that nugget score differences align with human preferences, reinforcing the potential of nugget-based metrics as reliable evaluators of model quality in head-to-head evaluations.
@@
Figure 4: Confusion matrix comparing human and nugget preferences. A threshold of 0.07 is applied to treat nugget preference scores as a tie.
Figure 4 presents a confusion matrix that visually compares the distribution of human preferences (rows) in Search Arena against “nugget preferences” (columns). For “nugget preference”, we use a threshold of 0.07, meaning that when the nugget score difference between the two model outputs falls within ±$\pm~{}0.07$, the comparison is considered a tie. This threshold was selected by sweeping values between 0.05 and 0.15 in increments of 0.01. A threshold of 0.07 results in nugget preferences that most closely reflect an equal distribution of modelA wins, modelB wins, and ties when the human preference is a tie (second row in Figure 4). The diagonal cells in this confusion matrix reveal the instances where nugget preferences align with human preferences. Conversely, off-diagonal cells illustrate the types and frequencies of disagreements between the human and nugget scores.
In particular, the nugget-based evaluation prefers modelA in 938 out of 1,714 (54.7%) of the battles where modelA wins the battle (first row in Figure 4). Similarly, modelB is preferred in 924 out of 1761 (52.5%) battles where it wins the battle (third row in Figure 4). We investigate the anti-diagonal corners where nugget and human preferences disagree–in the first two studies below. Then, we examine how the availability of URL contents affects nugget generation. Finally, we compare LLM-as-a-judge preference agreement with human preferences as an alternative to relying on the nugget-based preferences.
4.1 Query Classification Analysis
In this analysis, we use query classification to better understand the cases where nugget preferences and human preferences are not aligned. When the nugget scores and the human prefer opposite sides of a battle, we refer to this situation as a “preference inversion”, or simply inversion.
We suspect that inversions might vary across different types of queries. To investigate, we followed Rosset et al. [26] but used the newer GPT4.1 to rate each query on a scale of 0–10 across eight different categories. Then, we classify each query into its maximum scoring category or categories (allowing for ties). To further strengthen the category signals, we exclude queries with a maximum score less than seven from this classification. Raw distributions of the query ratings per category and sample queries from each class are available in Section A.2.
(a) Ambiguous
(b) Assumptive
(c) Multi-faceted